OpenStudy (anonymous):

Find a and b. The parabola y=ax^2+bx-3 has vertex (3,15).

myininaya (myininaya):

Have you tried using the vertex formula for your equation y=ax^2+bx+c (-b/(2a), f(-b/(2a))

myininaya (myininaya):

\[(3,15)=(\frac{-b}{2a},f(\frac{-b}{2a}))\]

OpenStudy (anonymous):

yes but I keep getting it wrong therefore would you please help with the answer?

myininaya (myininaya):

yep. just wondering if you had tried it.

myininaya (myininaya):

so we have two equations: \[3=\frac{-b}{2a} \] \[15=f(\frac{-b}{2a})\]

myininaya (myininaya):

\[3=\frac{-b}{2a} => 6a=-b => -6a=b\] Go back to your y=f(x) equation and do f(3) and replace the b with -6a.

myininaya (myininaya):

Show me what you have when you do this.

myininaya (myininaya):

\[f(x)=ax^2+bx-3\] Replace the x's with 3 and the b with (-6a) and then remember we can say f(-b/(2a)) or f(3)=15 since -b/(2a) is 3.

OpenStudy (anonymous):

so a= 2 and b= 3>

myininaya (myininaya):

that wouldn't work. 3 does not equal -3/(2*2) Did you try what I said? I was hoping you would just perform the one step I asked you to do.

OpenStudy (anonymous):

a = -3 and b = -15?

myininaya (myininaya):

We have \[f(x)=ax^2+bx-3 \] I asked you to replace the x's with 3's like this: \[f(3)=a(3)^2+b(3)-3\] So this means after simplifying we have \[f(3)=a(9)+b(3)-3\] Now I also asked you to replace the b with (-6a) like this \[f(3)=9a+(-6a)(3)-3\]

myininaya (myininaya):

now remember f(3)=15 so we have: \[9a+(-6a)(3)-3=15 \]

myininaya (myininaya):

Do you think you can solve this for a?

OpenStudy (anonymous):

no, i'm so confused .. I'm so sorry, this is literally the 60th math question i'm finishing up today.

myininaya (myininaya):

Can you tell me what confuses you? Like just the whole problem?

OpenStudy (anonymous):

the whole problem but i'll ask my teacher to give me a hand on monday.

myininaya (myininaya):

So you don't have any clue that the vertex of \[y=ax^2+bx+c \text{ is } (\frac{-b}{2a}, f(-\frac{b}{2a}))?\]

myininaya (myininaya):

Like do I need to convince you why that is? Or is it something else?

OpenStudy (anonymous):

no clue :(

myininaya (myininaya):

What? I'm confused. What do you mean? No clue what I'm asking? Or you don't know that is the vertex formula for anything written in this form y=ax^2+bx+c where a does not equal 0?

OpenStudy (anonymous):

oh the vertex , what should I do to get it?

myininaya (myininaya):

You already have the vertex. The vertex is (3,15).

myininaya (myininaya):

But we know that \[(3,15)=(\frac{-b}{2a}, f(\frac{-b}{2a}))\]

myininaya (myininaya):

since that is the vertex formula for y=ax^2+bx+c where a does not equal 0 and the vertex was actually given (3,15)

myininaya (myininaya):

So we want both of those x-coordinates to be equal and we want both of those y-coordinates to be equal

OpenStudy (anonymous):

oh yes, i meant the values of a and b

myininaya (myininaya):

\[3=\frac{-b}{2a} \text{ <-the x-'s } \text{ the y's -> } 15=f(-\frac{b}{2a})\] Or we could just say \[15=f(3) \text{ since } f(\frac{-b}{2a})=f(3) \]

myininaya (myininaya):

f(3)=15 means we have \[f(3)=a(3)^2+b(3)-3=15\] \[a(3)^2+b(3)-3=15 \]

myininaya (myininaya):

now remember you also have \[3=\frac{-b}{2a}\] Solve this for b.

myininaya (myininaya):

I'm trying to get you to get b by itself. b is being divided by (-2a) so multiply both sides by (-2a)

OpenStudy (anonymous):

b = -6a

myininaya (myininaya):

Right. now go to the equation I posted that had the other a and b and replace that b with (-6a)

myininaya (myininaya):

The equation right above the one I asked you to solve for b.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so what do i do?

myininaya (myininaya):

I'm still waiting for you to show me the rewrite of equation I wrote above the one you solved for b. You know I asked you to replace that b with (-6a) since b equals (-6a)

myininaya (myininaya):

Remember we have also \[a(3)^2+b(3)-3=15 \]

myininaya (myininaya):

I'm trying to get you to replace that b with (-6a) just put (-6a) instead of b that is all I'm asking.

myininaya (myininaya):